An ideal gas (Cp/Cv = γ) is taken through a process in which the pressure and the volume vary as p = aVb. Find the value of b for which the specific heat capacity in the process is zero.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. C base P/C base V = γ, C base P – C base V = R, C base V = r/γ -1, C base P = γR/γ – 1| Pdv = 1/b + 1(Rdt) ⇒ 0 = C base VdT + 1/b + 1(Rdt) ⇒ 1/b + 1 = -C base V/R ⇒ b + 1 = -R/C base V = -(C base P – C base V)/C base V = -γ + 1 ⇒ b = -γ